# RLTHF: Targeted Human Feedback for LLM Alignment

### v2

@@ Appendix B Iterative Alignment Improvement
Appendix C Experimental Setup
C.1 Data Preparation
C.1.1 Datasets
@@
C.1.2 Flipping human preferences
@@
To minimize this unfair penalty in our evaluation, and following prior work (Wang et al., 2024a), we first train an RM using the full set of original human annotations. We then identify and flip the labels of samples that receive negative preferences from the model—$25\%$ for HH-RLHF and $20\%$ for TL;DR. These flipped labels serve as the ground truth for all experiments.
@@
Preference Source for DPO HH-RLHF TL;DR
@@ C.2 Model Training
All training is done on a node of 8$\times$×A100 NVIDIA GPUs with DeepSpeed.
C.3 Baselines
@@
C.4 RLTHF-Specific Configurations
@@
  • Amplification Ratio: The default value of $\alpha$ is set to 4.
@@
Appendix D Obtaining Pair-wise Win Rate with AlpacaEval
@@
D.1 Adapted Prompt Template for HH-RLHF
@@
D.2 Adapted Prompt Template for TL;DR

### v3

@@ Appendix C Iterative Alignment Improvement
Appendix D Experimental Setup
D.1 Data Preparation
D.1.1 Datasets
@@
D.1.2 Flipping human preferences
@@
[…]
@@
Preference Source for DPO $H$-RLHF TL;DR
@@ D.2 Model Training
[…]
D.3 Baselines
@@
D.4 RLTHF-Specific Configurations
@@
  • Amplification Ratio: The default $\alpha$ values are 4, 4, 4, 2, and 1 for iterations 1–5, respectively, and 1 for all subsequent iterations.
@@
Appendix E Obtaining Pair-wise Win Rate with AlpacaEval
@@
E.1 Adapted Prompt Template for HH-RLHF
@@
E.2 Adapted Prompt Template for TL;DR
@@
Appendix F Cost Analysis
We take our experiments on HH-RLHF as a case study.
Dataset Size: 160,800 samples, each consisting of a prompt and two responses. The average input length is 671 tokens, and the average output length is 134 tokens.
Human Annotation Cost: Amazon Mechanical Turk (Amazon Web Services, 2025) suggested text classification pricing:
$0.012\times 3\text{ (labelers)}=\$0.036\text{ per sample}$
Note: Here the suggested pricing may be much lower than the actual cost. Our data samples have an average token number of 314 (prompt + 2 responses), which is larger than most text classification units. AMT’s labeling service providers typically list an hourly rate of \$6–7. According to human reading speed of 200–250 words per minute, the actual cost should be around \$0.13–0.18/sample/labeler, which is more than $10\times$ of the suggested pricing. In the following analysis, we still use the suggested pricing as a lower bound to provide a conservative estimate of RLTHF’s gain.
LLM Annotation Cost: Table 4 summarizes the estimated LLM annotation cost per sample for GPT-4o and GPT-4o mini, based on OpenAI’s token-based pricing (OpenAI, 2025) [^1] Pricing as of March 31, 2025..
Table 4: GPT-4o and GPT-4o mini Annotation Cost
Model Input Cost Input Tokens Output Cost Output Tokens Avg Cost
(\$ per 1M tokens) (\$ per 1M tokens) (\$ per sample)
GPT-4o 2.50 671 10.00 134 0.0030
GPT-4o mini 0.15 0.60 0.00018
[…]
Comparison: For computing, we only consider RM training and inference, as the downstream LLM training is the same for both full-human annotation and RLTHF. Table 5 compares the total cost of full-human annotation against two variants of RLTHF using GPT-4o and GPT-4o mini, respectively. The RLTHF setting assumes only 6% human annotation, 1/4 dataset shard for training, and 7 RM training iterations.
Table 5: Cost Comparison between Full-human Annotation and RLTHF
Solution Human Annotation (\$) LLM Annotation (\$) RM Train & Infer (\$) Total (\$)
Full-human $0.036\times 160{,}800=5788.8$ 5788.8
RLTHF (4o) $036\times 160{,}800\times 0.06=347.3 0.0$$0030\times\frac{160{,}800}{4}=120.6$ $32.77\times 2\times 7=458.8$ 926.7
RLTHF (4o mini) $36\times 160{,}800\times 0.06=347.3 0.000$$0018\times\frac{160{,}800}{4}=7.2$ $32.77\times 2\times 7=458.8$ 813.3
Even counting the extra LLM labeling and computing overhead, RLTHF can still reduce the overall cost by 84.0–86.0%. Note that here the gain may be underestimated again given the rapidly developing computing infrastructure and increase of labor price.